# Large Language Models Can Help Mitigate Barren Plateaus

### v1

@@ Abstract
In the era of noisy intermediate-scale quantum (NISQ) computing, Quantum Neural Networks (qnns) have emerged as a promising approach for various applications, yet their training is often hindered by barren plateaus (BPs), where gradient variance vanishes exponentially as the model size increases. To address this challenge, we propose a new Large Language Model (llm)-driven search framework, AdaInit, that iteratively searches for optimal initial parameters of qnns to maximize gradient variance and therefore mitigate BPs. Unlike conventional one-time initialization methods, AdaInit dynamically refines qnn’s initialization using llms with adaptive prompting. Theoretical analysis of the Expected Improvement (EI) proves a supremum for the search, ensuring this process can eventually identify the optimal initial parameter of the qnn. Extensive experiments across four public datasets demonstrate that AdaInit significantly enhances qnn’s trainability compared to classic initialization methods, validating its effectiveness in mitigating BPs.
@@ 1 Introduction
In recent years, there have been significant advancements in quantum computing, particularly with the advent of noisy intermediate-scale quantum (NISQ) devices Preskill (2018). Within this research landscape, quantum neural networks (qnns), which integrate quantum circuits with classical deep-learning layers, have been widely applied in various domains, such as quantum machine learning Stein et al. (2021); Liang et al. (2024), quantum physics Chen et al. (2022), and quantum hardware architecture Liang et al. (2023). However, recent studies reveal that the performance of qnns may be hindered due to gradient issues, such as barren plateaus (BPs) McClean et al. (2018), referring to a kind of gradient issue that the initialization of qnns might be trapped on a flattened landscape at the beginning of training. McClean et al. (2018) first systematically investigate BPs and affirm that the gradient variance will exponentially decrease as the model size increases when the qnns satisfy the assumption of the 2-design Haar distribution. Under this circumstance, most gradient-based approaches would fail. To better illustrate the BPs’ mitigation process, we present an example in Fig. 1.
@@
Figure 1: Example of BPs’ mitigation process. A flattened loss landscape ($1^{st}$ image), a.k.a. BPs, could be gradually recovered to the normal case ($3^{rd}$ image) by applying mitigation methods.
Numerous studies have been devoted to mitigating the barren plateau issues, whereas within these studies, initialization-based strategies have proven to be very effective by initializing qnns’ parameters with well-designed distributions Sack et al. (2022). However, most initialization-based strategies aim to mitigate BPs by one-time initialization with a well-designed data distribution, which may not be generalized to common data distribution. Initializing qnns’ parameters using deep-learning generative models could be a feasible solution since they can adaptively model data distribution on various datasets Friedrich and Maziero (2022). Within the category of generative models, large language models (llms) have demonstrated their superior performance in recent years Achiam et al. (2023); Dubey et al. (2024); Pham et al. (2024); Wang et al. (2024). Nonetheless, until now, leveraging the superior generative performance of llms to alleviate BPs is still under-explored.
To fill this research gap, we propose a new llm-driven framework, namely AdaInit, that can Adaptively generate Initial model parameters $\theta_{0}$ for qnns. After iterative generation, our framework obtains such an initial model parameter that can maximize the gradient variance for qnns’ training. Specifically, for each iteration, we estimate the posterior of $\theta_{0}$ using a generative model, such as a llm, given an adaptively improved prompt and a prior distribution as inputs. After posterior estimation, we train a qnn initialized with the generated $\theta_{0}$ and compute the gradient variance. We then evaluate the gradient variance by expected improvement (EI) and update the prompts if the EI is improved. Besides updating prompts, we store the corresponding $\theta_{0}$ and return the optimal one at the end. In this study, we theoretically analyze the submartingale property of EI and rigorously prove that the iterative search can eventually reach a supremum, which indicates that our framework can ultimately identify the optimal initial model parameters that maximize the gradient variance. Besides, we conduct extensive experiments to demonstrate the effectiveness of our proposed framework across four public datasets. The results reveal that our framework can maintain higher gradient variances against three classic initialization methods and two popular initialization-based strategies for mitigating BPs. Overall, we summarized our main primary contributions as follows:
@@
  • We propose a new llm-driven framework, AdaInit, for mitigating BPs. To the best of our knowledge, we first leverage llms to model qnns’ initial parameters for adaptively mitigating BPs.
@@
  • We theoretically analyze the submartingale property of expected improvement (EI) and rigorously prove the supremum of iterative search, providing theoretical validation for our search framework.
@@
  • Extensive experiments across four public datasets demonstrate that as the model size of qnns increases, our framework can maintain higher gradient variances against classic initialization methods.
2 Methodology
In this section, we first introduce the preliminary background and formally state the problem we aim to address in this study. Besides, we present our proposed framework in detail and conduct a theoretical analysis of the expected improvement (EI).
2.1 Preliminary Background
@@
Variational Quantum Circuits (vqcs) play a core role in quantum neural networks (qnns). Typical vqcs consist of a finite sequence of unitary gates $U(\bm{\theta})$ parameterized by $\bm{\theta}\in\mathbb{R}^{LNR}$, where $L$, $N$, and $R$ denote the number of layers, qubits, and rotation gates. $U(\bm{\theta})$ can be formulated as:
@@
where $U_{l}(\theta_{l})=e^{-i\theta_{l}V_{l}}$.
@@
where we denote $U_{-}\equiv\prod_{l=0}^{k-1}U_{l}(\theta_{l})W_{l}$ and $U_{+}\equiv\prod_{l=k}^{L}U_{l}(\theta_{l})W_{l}$. Also, $U(\bm{\theta})$ is sufficiently random s.t. both $U_{-}$ and $U_{+}$ (or either one) are independent and match the Haar distribution up to the second moment.
Barren Plateaus (BPs) are first investigated by McClean et al. (2018), who demonstrate that the gradient variance $\mathrm{Var}[\partial E]$ of qnns will exponentially decrease as the number of qubits $N$ increases when the random qnns match 2-design Haar distribution. This exponential pattern can be approximated as:
@@
By leveraging a generative AI (GenAI) model, such as an llm, as a Bayesian posterior estimator with adaptive prompting, we aim to iteratively identify the optimal qnn’s parameter $\bm{\theta^{*}_{0}}$, where a given qnn is initialized with $\bm{\theta^{*}_{0}}$, which can maximize gradient variance $\mathrm{Var}[\partial E]$ during training, thereby mitigating barren plateaus (BPs).
2.2 Our Proposed Framework
In this study, we introduce a new framework, AdaInit, designed to mitigate BP issues in qnns by leveraging generative AI (GenAI) models, particularly llms. Our key innovations can be described as follows. (i) First, unlike conventional one-time initialization strategies, we propose a generative approach that iteratively searches the optimal initial model parameters $\bm{\theta^{*}_{0}}\in\mathbb{R}^{LNR}$ that maximize the gradient variance $\mathrm{Var}[\partial E]$ of qnns, thereby mitigating BPs and improving qnns’ trainability. In each search iteration, we employ an llm as a Bayesian estimator to refine the posterior (candidate initial model parameters $\bm{\theta_{0}}$) through adaptive prompting. After posterior estimation, we train the qnn initialized with the generated $\bm{\theta_{0}}$ and further compute its $\mathrm{Var}[\partial E]$. The benefit of using llm as a posterior estimator is that the llm can incorporate diverse textual instructions via prompts and adaptively update the prompts based on feedback from the previous iteration. This adaptive refinement allows our framework to dynamically optimize the generation process. (ii) To validate the generation quality, we employ Expected Improvement (EI), $\Delta^{(t)}$, as a guiding metric for our search. Furthermore, we rigorously prove that the EI and its cumulative sum satisfy the properties of submartingale. Consequently, we theoretically establish their boundedness, thereby demonstrating that our proposed search framework will ultimately find the optimal initial model parameters for qnns.
@@ 2.2 Our Proposed Framework
Figure 2: Our proposed framework follows an iterative search process over $T$ iterations (gray area). In $t$-th iteration, we perform four sequential steps: (i) Generate $\bm{\theta}_{0}^{(t)}$ using a Gen AI model, $f(\cdot)$, (ii) Compute $\mathrm{Var}[\partial E^{(t)}]$ after qnn’s training, (iii) Calculate EI, $\Delta^{(t)}$, and (iv) Update prompts $x_{p}^{(t+1)}$, historical maximum gradient variance $S^{(t)}$, and optimal candidates $\bm{\theta}_{0}^{*}$ for next iteration. Dashed arrows indicate data flow and corresponding outputs in each step.
@@
Algorithm 1 Search for optimal initial model parameters for qnns.
0:  A GenAI model $f(\cdot)$, prompts $x_{p}$, a qnn $g(\cdot)$, the number of search iterations $T$.
1:  Initialize prompts $x_{p}$ and the GenAI model, $f(\cdot)$;
@@
2:  Create an empty list $\bm{\Theta}_{0}^{*}\leftarrow\varnothing$ to collect optimal candidates of initial model parameters for the qnn, $g(\cdot)$;
@@
7:     if $\Delta^{(t)}>\frac{1}{poly(N,L)T}$ then
@@
13:  return  $\bm{\Theta}_{0}^{*}$;
We present our framework workflow in Fig. 2 and further introduce details in Algo. 1. Given a GenAI model $f(\cdot)$, prompts $x_{p}$ for the $f(\cdot)$, a qnn $g(\cdot)$, and the number of search iterations $T$, we first initializes $f(\cdot)$, $x_{p}$ (line 1) and also creates an empty list $\varnothing$ for $\bm{\Theta}_{0}^{*}$ to collect optimal candidates of qnn’s initial model parameters (line 2). After initialization, we conduct $T$ iterations for searching (line 3). In each iteration, let’s say in the $t$-th iteration, we first employ $f(\cdot)$ with prompts $x_{p}^{(t)}$ and a prior distribution $P(\bm{\theta}_{0}^{(t)})$ to estimate the posterior distribution $P(\bm{\theta}_{0}^{(t)}|x_{p}^{(t)})$, which is the generated initial model parameter $\bm{\theta}_{0}^{(t)}$ for the qnn (line 4). After generation, we train the qnn $g(\bm{\theta}_{0}^{(t)})$ with certain training epochs and compute the gradient variance $\mathrm{Var}[\partial E^{(t)}]$, whose gradient is abbreviated from $\frac{\partial E(\bm{\theta}^{(t)})}{\partial\bm{\theta}^{(t)}}$, where $\bm{\theta}^{(t)}$ denotes the qnn’s model parameter in the $t$-th iteration (line 5). After computing the variance, we evaluate the improvement using the Expected Improvement (EI) metric, comparing the current gradient variance $\mathrm{Var}[\partial E^{(t)}]$ to the historical maximum gradient variance, which is the cumulative sum of EI when EI meets the following conditions (line 6). If the current EI, $\Delta^{(t)}$, is effectively improved, i.e., $\Delta^{(t)}>\frac{1}{poly(N,L)T}$, where $\frac{1}{poly(N,L)T}$ denotes a strictly positive lower bound on the gradient variance of an $N$-qubit, $L$-layer qnn for each search, in the absence of BPs (line 7), then we update the prompts for next iteration based on the current initial model parameters $\bm{\theta}_{0}^{(t)}$, the current gradient variance $\mathrm{Var}[\partial E^{(t)}]$, and the historical maximum gradient variance $S^{(t-1)}$ (line 8). After updating prompts, we update the historical maximum $S^{(t)}$ for the next iteration, where $S^{(t)}=S^{(t-1)}+\Delta^{(t)}=\mathrm{Var}[\partial E^{(t)}]$ (line 9) and further concatenate $\bm{\theta}_{0}^{(t)}$ to the optimal candidate list $\bm{\Theta}_{0}^{*}$ (line 10), which will be returned at the end (line 13). If so, the optimal initial model parameter $\bm{\theta}_{0}^{*}$ will be the last element in the candidate list.
@@
The search runs $T$ iterations. In each iteration, posterior estimation, which is linearly related to the output size of $\bm{\theta}_{0}$, takes $\mathcal{O}(|\bm{\theta}_{0}|)$ for a fixed-size qnn. Besides, training $g(\bm{\theta}_{0})$ with $T_{tr}$ epochs may take $\mathcal{O}(T_{tr}\cdot|\bm{\theta}_{0}|)$, where $T_{tr}$ denotes the number of training epochs for qnn. Combining $\bm{\theta}_{0}\in\mathbb{R}^{LNR}$, the total time complexity is $\mathcal{O}(T\cdot(L\cdot N\cdot R+T_{tr}\cdot L\cdot N\cdot R))\approx\mathcal{O}(T\cdot T_{tr}\cdot L\cdot N\cdot R)$. The space complexity primarily depends on the storage requirements. $\bm{\Theta}_{0}^{*}$ at most stores $T$ number of $\bm{\theta}_{0}$, which consumes $\mathcal{O}(T\cdot|\bm{\theta}_{0}|)$. The output of posterior estimation takes $\mathcal{O}(|\bm{\theta}_{0}|)$ space. Gradient variance and EI are scalars, which cost $\mathcal{O}(1)$ space. The prompts $x_{p}$ are iteratively updated and thus occupy $\mathcal{O}(|x_{p}|)$ space. Considering the size of $\bm{\theta}_{0}$, the total space complexity is $\mathcal{O}(T\cdot L\cdot N\cdot R+L\cdot N\cdot R+|x_{p}|)\approx\mathcal{O}(T\cdot L\cdot N\cdot R+|x_{p}|)$.
2.3 Theoretical Analysis of Expected Improvement (EI).
Before presenting the details of all necessary theoretical analysis, we would like to discuss how we can interpret these results. First, we formally define the Expected Improvement (EI) at each search iteration $t$ as $\Delta^{(t)}$ and its accumulative sum in the past iterations as $S^{(t-1)}$ in Def. 1. Besides, we assume that the maximum possible gradient $\partial E_{max}$ during qnn’s training is bounded by a positive constant $B_{\partial E}$, which is practical in real-world simulation. Next, we establish an upper bound for EI through Lem. 1 and Lem. 2. These results indicate that $S^{(t)}$ is $L^{1}$-bounded and integrable for each $t$. Building upon these lemmas, we investigate the submartingale property of $\Delta^{(t)}$ and rigorously prove in Lem. 3 that $S^{(t)}$ is submartingale. This insight is crucial as it provides a theoretical basis to analyze the convergence of our proposed search framework. Finally, leveraging the convergence of submartingales and the monotonicity of $S^{(t)}$, we establish in Lem. 4 that $S^{(t)}$ has a supremum, which indicates that our proposed search framework can eventually identify the optimal initial model parameters that maximize the gradient variance of qnns in optimization. Due to the page limit, we provide rigorous proof in the Appendix.
@@
For $\forall$ $t\in\mathbb{Z}^{+}$, the Expected Improvement (EI) in the $t$-th search iteration is defined as:

### v2

# Mitigating Barren Plateaus in Quantum Neural Networks via an AI-Driven Submartingale-Based Framework

@@ Abstract
In the era of noisy intermediate-scale quantum (NISQ) computing, Quantum Neural Networks (qnns) have emerged as a promising approach for various applications, yet their training is often hindered by barren plateaus (BPs), where gradient variance vanishes exponentially in terms of the qubit size. Most existing initialization-based mitigation strategies rely heavily on pre-designed static parameter distributions, thereby lacking adaptability to diverse model sizes or data conditions. To address these limitations, we propose AdaInit, a foundational framework that leverages generative models with the submartingale property to iteratively synthesize initial parameters for qnns that yield non-negligible gradient variance, thereby mitigating BPs. Unlike conventional one-shot initialization methods, AdaInit adaptively explores the parameter space by incorporating dataset characteristics and gradient feedback, with theoretical guarantees of convergence to finding a set of effective initial parameters for qnns. We provide rigorous theoretical analyses of the submartingale-based process and empirically validate that AdaInit consistently outperforms existing initialization methods in maintaining higher gradient variance across various qnn scales. We believe this work may initiate a new avenue to mitigate BPs.
@@ 1 Introduction
Figure 1: Example of BPs’ mitigation process. A plateau-dominated loss landscape ($1^{st}$ image), a.k.a. BPs, could be gradually recovered to the normal case ($3^{rd}$ image) after mitigation.
In recent years, there have been significant advancements in quantum computing, particularly with the advent of noisy intermediate-scale quantum (NISQ) devices Preskill (2018). Within this research landscape, quantum neural networks (qnns), which integrate quantum circuits with classical deep-learning layers, have been widely applied in various domains, such as quantum machine learning Stein et al. (2021), quantum chemistry and materials modeling Kandala et al. (2017), and combinatorial optimization Farhi et al. (2014). However, recent studies Ortiz Marrero et al. (2021); Cerezo et al. (2021) reveal that the performance of qnns may be hindered due to gradient issues, such as barren plateaus (BPs), referring to a kind of gradient issue that the initialization of qnns might be trapped on a flattened landscape at the beginning of training. McClean et al. (2018) first systematically investigate BPs and affirm that the gradient variance will exponentially decrease when the qnns satisfy the assumption of the 2-design Haar distribution. Under these circumstances, most gradient-based training approaches would fail. To better illustrate the BPs’ mitigation process, we present an example in Fig. 1.
Numerous studies have been devoted to mitigating the BPs, among which initialization-based strategies have proven particularly effective by using well-designed distributions to initialize qnns’ parameters Sack et al. (2022). However, most existing initialization-based strategies, such as GaInit Zhang et al. (2022) or BeInit Kulshrestha & Safro (2022), rely on a one-shot generation of static initial parameters, which suffer from two critical limitations: (i) they often depend on idealized distribution assumptions, and (ii) they lack adaptability when scaling to various model sizes or data conditions. These limitations are particularly problematic where BPs hinder effective qnns’ training.
To address the limitations, we propose AdaInit, an iterative framework that combines generative modeling with the submartingale property. We prove its convergence, demonstrating that AdaInit Adaptively generates effective Initial parameters for qnns. These parameters ensure non-negligible gradient variance at the beginning of training, thereby mitigating BPs. Beyond the theoretical guarantee, AdaInit intuitively can help qnns locate initial parameters in the “non-flat” loss landscape, which prevents training from being trapped from the start. Our method is grounded by two key ideas. First, instead of statically pre-designing the initialization distribution, we leverage generative models, such as large language models (llms), to synthesize candidate parameters based on dataset descriptions and prior gradient feedback. This allows the initializer to actively explore the parameter space and adaptively refine the candidate. Second, by modeling the iterative process as a submartingale, we provide a theoretical guarantee that this process will almost surely converge within a finite number of iterations to initial parameters that yield non-negligible gradient variances. We model such a process because the submartingale property can depict the trend of expected improvement in gradient variance and guarantee convergence in finite time. Besides theoretical analysis, we conduct extensive experiments to validate the effectiveness of our proposed framework across various model scales. The results reveal that our framework can maintain higher gradient variances against three classic initialization methods and two popular initialization-based strategies for mitigating BPs. Overall, our primary contributions can be summarized as:
@@
  • We propose a new artificial intelligence (AI)-driven submartingale-based framework, AdaInit, for mitigating BPs. To the best of our knowledge, we open a new avenue to leverage llms with submartingale property to model qnns’ initial parameters for mitigating BPs.
@@
  • We theoretically analyze the submartingale property of the iterative process and rigorously prove its supremum and expected hitting time.
@@
  • Extensive experiments across various model scales demonstrate that as the model size of qnns increases, our framework can maintain higher gradient variances against classic initialization methods.
2 Related Work
McClean et al. (2018) first investigated barren plateau (BP) phenomena and demonstrated that under the assumption of the 2-design Haar distribution, gradient variance in qnns will exponentially decrease to zero at the beginning of training as the model size increases. In recent years, enormous studies have been devoted to mitigating BP issues in qnnQi et al. (2023). Cunningham & Zhuang (2024) categorize most existing studies into the following five groups. (i) Initialization-based strategies initialize model parameters with various well-designed distributions in the initialization stage Grant et al. (2019); Sack et al. (2022); Mele et al. (2022); Grimsley et al. (2023); Liu et al. (2023); Park & Killoran (2024); Kashif et al. (2024); Shang & Shi (2025). (ii) Optimization-based strategies address BP issues and further enhance trainability during optimization Ostaszewski et al. (2021); Suzuki et al. (2021); Heyraud et al. (2023); Liu et al. (2024); Sannia et al. (2024). (iii) Model-based strategies attempt to mitigate BPs by proposing new model architectures Li et al. (2021); Bharti & Haug (2021); Du et al. (2022); Selvarajan et al. (2023); Tüysüz et al. (2023); Kashif & Al-Kuwari (2024). (iv) To address both BPs and saddle points, Zhuang et al. (2024) regularize qnns’ model parameters via Bayesian approaches. (v) Rappaport et al. (2023) measure the BP phenomenon via various informative metrics.
3 Preliminaries
In this section, we first introduce the preliminary background about the basics of quantum computing and barren plateaus, and then present the necessary tools from probability theory.
Quantum Basics. A quantum state can be seen as a unit vector $|\psi\rangle$ in a complex Hilbert space $\mathcal{H}^{m}$, satisfying the normalization condition $\langle\psi|\psi\rangle=1$. We use the Dirac bra-ket notation, where ket $|\psi\rangle$, bra $\langle\psi|$ denote a column vector in $\mathbb{C}^{m}$ and its Hermitian conjugate (conjugate transpose), respectively. Any $|\psi\rangle$ can be written as a linear combination of computational basis states, $|\psi\rangle=\sum_{i=1}^{m}c_{i}|i\rangle$, where $c_{i}\in\mathbb{C}$ are called the amplitudes of the basis states $|i\rangle$. Given two states $|\psi\rangle\in\mathcal{H}^{m}$ and $|\phi\rangle\in\mathcal{H}^{n}$, their inner product can be denoted by $\langle\psi|\phi\rangle\triangleq\sum_{i}\psi^{\dagger}_{i}\phi_{i}$, whereas their tensor product can be denoted by $|\psi\rangle\otimes|\phi\rangle\in\mathcal{H}^{m\times n}$. If we measure state $|\psi\rangle=\sum_{i=1}^{m}c_{i}|i\rangle$ on a computational basis, we will obtain $i$ with probability $|c_{i}|^{2}$ and the state will collapse into $|i\rangle$ after measurement for all $i$. A quantum algorithm works by applying a sequence of unitary operators to an initial state.
@@ 3 Preliminaries
Variational Quantum Circuits (vqcs) play a core role in quantum neural networks (qnns). Typical vqcs consist of a finite sequence of unitary gates $U(\bm{\theta})$ parameterized by $\bm{\theta}\in\mathbb{R}^{LNR}$, where $L$, NN, and $R$ denote the number of layers, qubits, and rotation gates. $U(\bm{\theta})$ can be formulated as:
@@
where $U_{l}(\theta_{l})=e^{-i\theta_{l}V_{l}}$, $V_{l}$ is a Hermitian operator.
@@
where we denote $U_{-}\equiv\prod_{l=0}^{k-1}U_{l}(\theta_{l})$ and $U_{+}\equiv\prod_{l=k}^{L}U_{l}(\theta_{l})$. Also, $U(\bm{\theta})$ is sufficiently random s.t. both $U_{-}$ and $U_{+}$ (or either one) are independent and match the Haar distribution up to the second moment.
Barren Plateaus (BPs) are first investigated by McClean et al. (2018), who demonstrate that the gradient variance $\mathrm{Var}[\partial E]$ of qnns at the beginning of training will exponentially decrease as the number of qubits $N$ increases when the random qnns match 2-design Haar distribution. This exponential pattern can be approximated as:
@@
By leveraging a generative AI (GenAI) model, such as an llm, we refine posterior with adaptive prompting, i.e., iteratively generate effective initial parameters $\bm{\theta^{*}_{0}}$ for a qnn that yields non-negligible gradient variance $\mathrm{Var}[\partial E]$, thereby mitigating barren plateaus (BPs).
Tools from Probability Theory. Below, we review the definition of martingale (submartingale), along with key tools relevant to our work. We adapt the descriptions from Williams (1991); Freeman & Stephenson (2025).
Definition 1 (Martingale, Williams (1991)).
Let $\{M^{(t)}\}_{t\geq 1}$ be a stochastic process w.r.t. a filtration $\{\mathcal{F}^{(t)}\}_{t\geq 1}$ on a probability space $(\Omega,\mathcal{F},P)$. The process $\{M^{(t)}\}$ is called a martingale if (i) $\{M^{(t)}\}$ is adapted, (ii) $\mathbb{E}[|M^{(t)}|]<\infty$, for $\forall\ t\in\mathbb{Z}^{+}$, (iii) $\mathbb{E}[M^{(t+1)}\mid\mathcal{F}^{(t)}]=M^{(t)}$, almost surely for $\forall\ t\in\mathbb{Z}^{+}$.
If (iii) is replace by $\mathbb{E}[M^{(t+1)}\mid\mathcal{F}^{(t)}]\geq M^{(t)}$ almost surely, we say $\{M^{(t)}\}$ is a submartingale.
Theorem 1 (Doob’s Forward Convergence Theorem, Williams (1991)).
Let $\{M^{(t)}\}_{t\geq 1}$ be an $L^{1}$-bounded submartingale (in Def. 1). Then, almost surely, the limit $M^{(\infty)}=\lim_{t\to\infty}M^{(t)}$ exists and is finite.
Theorem 2 (Doob’s Optional Stopping Theorem, Williams (1991)).
Let $\{M^{(t)}\}_{t\geq 1}$ be a submartingale (in Def. 1) and let $\tau$ be a stopping time. Then $\mathbb{E}[M^{(\tau)}]\geq\mathbb{E}[M^{(0)}]$ if any one of the following conditions hold: (i) $\tau$ is bounded, (ii) $P[\tau<\infty]=1$ and $\{M^{(t)}\}$ is bounded for $\forall\ t\in\mathbb{Z}^{+}$, (iii) $\mathbb{E}[\tau]<\infty$ and $|M^{(t)}-M^{(t-1)}|$ is bounded for $\forall\ t\in\mathbb{Z}^{+}$.
Theorem 3 (Dominated Convergence Theorem, Williams (1991)).
Let $^{(t)}, MM$ be random variables s.t. $M^{(t)}\to M$ almost surely. There exists a random variable $Y\in L^{1}$ s.t. for $\forall\ t\in\mathbb{Z}^{+}$, $|M^{(t)}|<Y$ almost surely, then $\mathbb{E}[M^{(t)}]\to\mathbb{E}[M]$ as $t\to\infty$.
Lemma 1 (Minimum of Stopping Times, Freeman & Stephenson (2025)).
Let $\tau_{1}$ and $\tau_{2}$ be stopping times w.r.t. a filtration $\{\mathcal{F}^{(t)}\}$. Then $\tau_{1}\wedge\tau_{2}=\min(\tau_{1},\tau_{2})$ is also a $\{\mathcal{F}^{(t)}\}$ stopping time.
4 Our Proposed Framework
@@
Figure 2: Our proposed framework follows an iterative process over $T$ iterations (gray area). In $t$-th iteration, we perform four sequential steps: (i) Generate $\bm{\theta}_{0}^{(t)}$ using a Gen AI model, $f(\cdot)$, (ii) Compute $\mathrm{Var}[\partial E^{(t)}]$ after qnn’s training, (iii) Calculate EI, $\Delta^{(t)}$, and (iv) Update prompts $x_{p}^{(t+1)}$, historical maximum gradient variance $S^{(t)}$, and effective candidates $\bm{\theta}_{0}^{*}$ for next iteration. Dashed arrows indicate data flow and corresponding outputs in each step.
In this study, we introduce a new framework, AdaInit, designed to mitigate BP issues in qnns by leveraging generative AI (GenAI) models, particularly llms. Our key innovations can be described as follows. (i) First, unlike conventional one-shot initialization strategies, we propose a generative approach that iteratively generates effective initial model parameters $\bm{\theta^{*}_{0}}\in\mathbb{R}^{LNR}$ for qnns that yield non-negligible gradient variance $\mathrm{Var}[\partial E]$, thereby mitigating BPs. In each iteration, we employ an llm to refine the posterior (candidate initial model parameters $\bm{\theta_{0}}$) through adaptive prompting. After the posterior refinement, we train the qnn initialized with the generated $\bm{\theta_{0}}$ and further compute its $\mathrm{Var}[\partial E]$ in the early stage of training for monitoring BPs. The benefit of using llm to refine the posterior is that the llm can incorporate diverse textual instructions via prompts and adaptively update the prompts based on feedback from the previous iteration. This adaptive refinement allows our framework to dynamically optimize the generation process. (ii) To validate the generation quality, we employ Expected Improvement (EI), $\Delta^{(t)}$, as a guiding metric for the iterative process. Furthermore, we rigorously prove that the process satisfies the properties of a submartingale. Consequently, we theoretically establish the boundedness, thereby demonstrating that our proposed framework will ultimately find effective initial model parameters for qnnsin a finite step.
@@
Algorithm 1 Iteratively generate effective initial parameters for qnns using a generative model.
0: A GenAI model $f(\cdot)$, prompts $x_{p}$, a qnn $g(\cdot)$, the number of iterations $T$, a parameter $K$.
1: Initialize prompts $x_{p}$ and the GenAI model $f(\cdot)$;
@@
2: Create an empty list $\bm{\Theta}_{0}^{*}\leftarrow\varnothing$ to collect effective candidates of initial model parameters for the qnn, $g(\cdot)$;
@@
7:  if $\Delta^{(t)}\geq\frac{1}{poly(N,L)K}$ then
@@
13:return $\bm{\Theta}_{0}^{*}$.
We present our framework workflow in Fig. 2 and further introduce details in Algo. 1. Given a GenAI model $f(\cdot)$, prompts $x_{p}$ for the $f(\cdot)$, a qnn $g(\cdot)$, and the number of iterations $T$, we first initialize $f(\cdot)$, $x_{p}$ (line 1) and also create an empty list $\varnothing$ for $\bm{\Theta}_{0}^{*}$ to collect candidates of qnn’s initial model parameters (line 2). After initialization, we run $T$ iterations for generation (line 3). In each iteration, let’s say in the $t$-th iteration, we first employ $f(\cdot)$ with prompts $x_{p}^{(t)}$ and a prior distribution $P(\bm{\theta}_{0}^{(t)})$ to refine the posterior distribution $P(\bm{\theta}_{0}^{(t)}|x_{p}^{(t)})$, which is the generated initial model parameter $\bm{\theta}_{0}^{(t)}$ for the qnn (line 4). After generation, we train the qnn $g(\bm{\theta}_{0}^{(t)})$ with certain training epochs and compute the gradient variance $\mathrm{Var}[\partial E^{(t)}]$, whose gradient is abbreviated from $\frac{\partial E(\bm{\theta}^{(t)})}{\partial\bm{\theta}^{(t)}}$, where $\bm{\theta}^{(t)}$ denotes the qnn’s model parameter in the $t$-th iteration (line 5). After computing the variance, we evaluate the improvement using the Expected Improvement (EI) metric, comparing the current gradient variance $\mathrm{Var}[\partial E^{(t)}]$ to the historical maximum gradient variance, which is the cumulative sum of EI when EI meets the following conditions (line 6). If the current EI, $\Delta^{(t)}$, is effectively improved, i.e., $\Delta^{(t)}\geq\nicefrac{{1}}{{(poly(N,L)K)}}$, where $\nicefrac{{1}}{{(poly(N,L)K)}}$ (with a parameter $K$ be determined later) denotes a strictly positive lower bound on the gradient variance of an $N$-qubit, $L$-layer qnn for each iteration, in the absence of BPs (line 7), then we update the prompts for the next iteration based on the current initial model parameters $\bm{\theta}_{0}^{(t)}$, the current gradient variance $\mathrm{Var}[\partial E^{(t)}]$, and the historical maximum gradient variance $S^{(t-1)}$ (line 8). After updating prompts, we update the historical maximum $S^{(t)}$ for the next iteration, where (if $\Delta^{(t)}\geq\nicefrac{{1}}{{(poly(N,L)K)}}$) $S^{(t)}=S^{(t-1)}+\Delta^{(t)}=\mathrm{Var}[\partial E^{(t)}]$ (line 9) and further concatenate $\bm{\theta}_{0}^{(t)}$ to the candidate list $\bm{\Theta}_{0}^{*}$ (line 10), which will be returned at the end (line 13). If so, the most effective initial model parameter $\bm{\theta}_{0}^{*}$ will be the last element in the candidate list. We can see that the input parameter $K$ is somewhat related to a wanted increment in $\mathrm{Var}[\partial E^{(t)}]$, which is further linked to the desired underlying property provided by the GenAI model. This connection is explicitly shown in the theoretical analysis below.
@@
Our framework runs $T$ iterations. In each iteration, posterior refinement, which is linearly related to the output size of $\bm{\theta}_{0}$, takes $\mathcal{O}(|\bm{\theta}_{0}|)$ for a fixed-size qnn. Besides, training $g(\bm{\theta}_{0})$ with $T_{tr}$ epochs may take $\mathcal{O}(T_{tr}\cdot|\bm{\theta}_{0}|)$, where $T_{tr}$ denotes the number of training epochs for qnn. Combining $\bm{\theta}_{0}\in\mathbb{R}^{LNR}$, the total time complexity is $\mathcal{O}(T\cdot(L\cdot N\cdot R+T_{tr}\cdot L\cdot N\cdot R))\approx\mathcal{O}(T\cdot T_{tr}\cdot L\cdot N\cdot R)$. The space complexity primarily depends on the storage requirements. $\bm{\Theta}_{0}^{*}$ at most stores $T$ number of $\bm{\theta}_{0}$, which consumes $\mathcal{O}(T\cdot|\bm{\theta}_{0}|)$. The output of posterior refinement takes $\mathcal{O}(|\bm{\theta}_{0}|)$ space. Gradient variance and EI are scalars, which cost $\mathcal{O}(1)$ space. The prompts $x_{p}$ are iteratively updated and thus occupy $\mathcal{O}(|x_{p}|)$ space. Considering the size of $\bm{\theta}_{0}$, the total space complexity is $\mathcal{O}(T\cdot L\cdot N\cdot R+L\cdot N\cdot R+|x_{p}|)\approx\mathcal{O}(T\cdot L\cdot N\cdot R+|x_{p}|)$.
Theoretical analysis of our framework.
We first present some necessary results and further discuss how those results can be interpreted. Full Proofs can be found in the Appendix A.
First, we define the Expected Improvement (EI) at each iteration $t$ as $\Delta^{(t)}$ and its cumulative sum in the past iterations as $S^{(t-1)}$ in Def. 2. Besides, we assume that the maximum possible gradient $\partial E_{max}$ during qnn’s training is bounded by a positive constant $B_{\partial E}$, which is practical in real-world simulation. Next, we establish an upper bound for EI through Lem. 2 and Lem. 3.
@@
For $\forall$ $t\in\mathbb{Z}^{+}$, the Expected Improvement (EI) in the $t$-th iteration is defined as: